The surface areas of two similar solids are 311 ft^2 and 1,037 ft^2. The volume of the larger solid is 1755 ft^3. What is the volume of the smaller solid? Also, the volume of two similar solids are 729m^3 and 125 m^3. The surface area of the larger solid is 324m^3. What is the surface area of the smaller solid?
Answer by Alan3354(30993)

The surface areas of two similar solids are 311 ft^2 and 1,037 ft^2. The volume of the larger solid is 1755 ft^3. What is the volume of the smaller solid?
-------------------
The ratio of area is the square of the ratio of any linear dimension.
The ratio of volume is the cube of the ratio of any linear dimension.
Linear ratio = sqrt(311/1037)
Volume ratio = (sqrt(311/1037))^3
Smaller volume = 1755*(sqrt(311/1037))^3
SV = 288.24 ft^3
-------------------
Also, the volume of two similar solids are 729m^3 and 125 m^3. The surface area of the larger solid is 324m^3. What is the surface area of the smaller solid?
Volume ratio = 125/729
Linear ratio = cube root of volume ratio = 5/9
Smaller area = 324*(5/9)^2 = 324*25/81 = 100 m^2 (you have m^3 for the area)